OpenStudy (anonymous):

find the consecutive integers such that the sum of the first two integers is 24 more than the third integer.

Parth (parthkohli):

\[\large x + 1, x+2, x + 3 \]These are your required numbers. Now, we have an equation:\[\large (x + 1) + (x + 2) = 24 + (x + 3)\]

OpenStudy (anonymous):

so now do i foil the equation?

Parth (parthkohli):

No, you're NOT multiplying!

OpenStudy (anonymous):

so am i combining like terms?

OpenStudy (anonymous):

let (x+1),(x+2),(x+3) be those numbers now according to ur question we get (x+1)+(x+2)=24+(x+3) solution 2x + 3 = x +27 combining like terms 2x - x = 27 - 3 x = 24 now ur numbers are 24,25,26